I am struggling with having to rearrange formula to make x the subject. For example y=2(x+1)
Also how do I solve the following equations to find a numerical value for x
example 2x+3=x+8 or 4(x+2)=20
Thankyou

Answer by Greenfinch(383)

Start with y=2x + 1
Move the 1 y - 1 = 2x
Divide by two (y-1)/2 = x or x= 1/2 x (y-1)
2x +3 = x + 8
Take x from each side x + 3 = 8
Take 3 from each side x = 5
And 4(x + 2) = 20
Remove bracket 4x + 8 = 20
Take 8 from each side 4x = 12
Divide by four x = 3